OpenStudy (anonymous):

find positive integers a,b such that : (\sqrt[3]{a}+\sqrt[3]{b}-1)^{^{2}}=49 + 20 \sqrt[3]6{}

OpenStudy (anonymous):

edit the question plz...

OpenStudy (anonymous):

\(\large (\sqrt[3]{a}+\sqrt[3]{b}-1)^{^{2}}=49 + 20 \sqrt[3]6{}\) ?

OpenStudy (anonymous):

yes, that is right :d

OpenStudy (anonymous):

well how about the question itself? finding all pairs of (a,b) or just one pair

OpenStudy (anonymous):

find all pairs :d

OpenStudy (anonymous):

lol.....very hard problem...

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

\[49+20\sqrt [3]{6}= (\sqrt [3]{48}+\sqrt [3]{288}-1)^{2}\] so \[(\sqrt [3]{a}+\sqrt [3]{b}-1)^{2}= (\sqrt [3]{48}+\sqrt [3]{288}-1)^{2}\] and \[\sqrt [3]{a}+\sqrt [3]{b}=\sqrt [3]{48}+\sqrt [3]{288}\]

OpenStudy (anonymous):

but having other pairs (a,b) ?

OpenStudy (anonymous):

i have to think again....but i guess there is one pair...(a,b)=(48,288)

OpenStudy (anonymous):

yes :d